2.2. Consider the following scenario. Initially the economy has 90 million people working, 10 

million people unemployed, and 20 million people not in the labor force. Then prospects for the 

economy improve. Five million people who previously were not in the labor force now join the 10 

million previously unemployed in looking for work. For now, the economy remains with 90 

million workers. What happens to the unemployment rate


1
Expert's answer
2021-08-16T08:49:03-0400

"Labour\space force=employed +unemployed\\=90+10\\=100\space million"

"Unemployment \space rate=\frac{unemployed}{labor \space force}×100\%\\\frac{10}{100}×100\%\\=10\%"


Five million people join 10 million previously unemployed people

New unemployed"=10+5=15\space million"

Unemployment rate"=\frac{15}{100}×100\%"

"=15\%"

Unemployment rate increases from 10% to 15%